Please help w/pre-test for final.
Find f^-1 (x), if it exists, for f(x) = 6x+5
Find the total amount of money in an account if $8000 was deposited 6 years ago at 2% interest compounded semi-annually.

Answer by stanbon(57407)

Find f^-1 (x), if it exists, for f(x) = 6x+5
You have y = 6x+5
Interchange x and y to get:
x = 6y + 5
Solve for y:
y = (1/6)x -(5/6)
--------------------------
Find the total amount of money in an account if $8000 was deposited 6 years ago at 12% interest compounded semi-annually.
A(6) = 8000(1+0.12/2)^(2*6)
A(6) = 8000(1.06)^12
A(6) = 8000*2.012196
A(6) = $16,097.57
========================
Cheers,
Stan H.